# Analytical Phasor-Based Fault Location Enhancement for Wind Farm Collector Networks

### v1

@@ Abstract
The increasing integration of Inverter-Based Resources (IBRs) is reshaping fault current characteristics, presenting significant challenges to traditional protection and fault location methods. This paper addresses a key limitation in fault location within wind farm collector networks, i.e., one-terminal phasor-based methods become inaccurate when IBRs are electrically located downstream from the fault. In such cases, the voltage drop caused by IBR fault current injections is not captured by the Intelligent Electronic Device, resulting in a systematic overestimation of fault distance. To mitigate this issue, a general compensation framework was proposed by augmenting classical loop formulations with a distance-dependent voltage correction term. The methodology was derived analytically using a sequence-domain representation and generalized to multiple fault types through a unified notation. It maintains the simplicity and interpretability of conventional approaches and can be implemented using only local measurements. The method was evaluated through EMT simulations in PSCAD using a realistic wind farm model. Results show significant improvements in location accuracy, with average and maximum errors notably reduced, especially for ground-involved faults where reductions exceed 90%. Furthermore, the compensation eliminates sensitivity to wind penetration levels and ensures uniform performance across feeders, positioning the method as a practical solution for modern renewable-dominated grids.
@@ I Introduction
Most fault-location categorized into four groups: phasor-based, voltage-sag, traveling-wave, and data-driven. Among these, phasor-based methods are often preferred because they are straightforward to implement, interpretable, and compatible with existing measurement infrastructures [DAS2014].
Within phasor-based methods, single-ended schemes use only local measurements to estimate the fault distance. Common methods include the impedance method [ZIEGLER2011], the reactance method [CAPAR2014], and Takagi-type methods [TAKAGI1982, DAS2014, SEL2018]. Two-ended schemes extend this approach by using synchronized measurements at both line terminals and, therefore, require a communications link. The main studies include [GIRGIS1992], [JJ1990], [PR2011], [HE2011]. Despite this variety, phasor-based methodologies are grounded in standard short-circuit assumptions associated with synchronous-machine behavior.
@@
In transmission systems with substantial IBR penetration, two complementary directions emerged: (i) benchmarking-driven adaptations of classical phasor/impedance methods; and (ii) sequence-aware single-ended formulations tailored to converter behavior. Along the first direction, the authors of [Davi2023IBRMultiMethod] benchmarked ten one- and two-terminal impedance-based schemes across tens of thousands of PSCAD cases spanning diverse IBR controls and short-circuit levels, and proposed a multi-method selector that reduced location error in the absence of communications. Along the second, [Mobashsher2024LocalIBR] derived a sequence-network model for IBRs showing that the apparent negative-sequence impedance became effectively infinite when grid codes did not mandate negative-sequence injection, and designed a single-ended locator that employed negative- and zero-sequence polarizing currents, incorporated remote-infeed and line-charging effects, and maintained accuracy across wide ranges of fault resistance and grid-code settings.
@@
Motivated by these gaps, this paper focuses on single-ended, phasor-based fault location for feeder lines within wind-farm collector systems. Specifically, it proposes a simple and transparent modification to the classical single-ended loop equation to account for the voltage drop induced by IBR currents. The resulting formulation preserves the simplicity and interpretability of classical phasor-based locators while improving accuracy in inverter-dominated networks.
@@ III-A Single-Line-to-Ground Fault Derivation
Together with the positive-sequence relation in Equation (9), these expressions fully specify the sequence-domain fault voltages ($V_{F}^{(1)},V_{F}^{(2)},V_{F}^{(0)}$). For other fault types, the per-sequence forms remain valid, i.e., what changed was the interconnection of the sequence networks at the fault. For instance, in an SLG fault the three sequence networks were connected in series, whereas in a line-to-line fault the positive- and negative-sequence networks were in series and the zero-sequence path was open.
@@ III-B Extension to Other Fault Types
Because the compensation term $V_{\text{comp}}(d)$ depends explicitly on the unknown distance $d$, its evaluation requires an iterative solution procedure. The estimation begins with an initial distance obtained from (16) by setting $V_{\text{comp}}(d)=0$. This preliminary value is then substituted into (17) to update the compensation term, yielding a revised value of $V_{\text{comp}}(d)$. The updated compensation is subsequently reintroduced into (16) to compute a refined estimate of $d$. This fixed-point iteration continues until successive distance estimates differ by less than a prescribed tolerance, ensuring convergence.
@@ III-C Practical Considerations and Iterative Estimation Procedure
Assuming that: (i) all inverter generations are equal during and previous the fault; and (ii) their positive-sequence injections maintain magnitude and phase angle close to pre-fault values, those assumptions can lead to an approximated form of Equation (17) as shown next:
@@ IV Test System
All WTGs were modeled using full-converter technology with a grid-following (GFL) control architecture [TREMBLAY2013, MILLER2003]. The GFL inverter synchronizes with the grid via a synchronous-reference-frame PLL [Se2000] and regulates the current through cascaded control loops. The inner current-control loop operates in the $dq$ frame and enforces the commanded current using PI controllers tuned via internal model control [Harne1998]. The outer loop manages active and reactive power set-points, with a default operation at a unity power factor during steady-state. The current magnitude was constrained to 1.1 $1\text{\,}\mathrm{p}\,\mathrm{u}\,$, and fault ride-through behavior was incorporated such that when the PCC voltage drops below $0.85\text{\,}\mathrm{p}\,\mathrm{u}\,$, the control system injects positive-sequence reactive current [Dadjo2023], as mandated by the Brazilian grid code [ONS2023].
@@ IV-A Monte Carlo Generation of Wind-Penetration Scenarios
The accuracy of the proposed fault-location methodology was influenced not only by the fault location and resistance, but also by the spatial distribution of wind-power injections along the collector system. The position of each wind turbine and its local operating point affect the pre-fault current sharing and, consequently, the short-circuit level seen at the relay. In practical installations, however, only the aggregated measurements at the IED bus are typically available, and the individual turbine currents are not recorded. In this work, each IBR contribution was therefore approximated as the pre-fault current at the IED divided by the number of WTGs (see Section III-C), implicitly assuming identical power output among turbines. Any dispersion in turbine output around this average may introduce additional location errors in the proposed methodology. To obtain a realistic and systematic assessment of this sensitivity, a Monte Carlo model was adopted to generate a large set of penetration scenarios that explicitly accounts for turbine-to-turbine variability.
@@
In this sense, across scenarios, the wind farm operates between $0\%$ and $100\%$ of rated power. The deviations were independently drawn from a symmetric uniform law as:
@@
Fig. 7 shows the evolution of $\varepsilon_{99}(n)$ as a function of the number of scenarios, together with the convergence threshold. As the number of simulations increases, the nearest-neighbor distance decreases approximately monotonically, indicating progressively finer resolution of the short-circuit current distribution. The resolution criterion (LABEL:eq:mc_convergence_criterion) was satisfied only when the Monte Carlo set approaches its final size, implying that additional simulations beyond this point would not introduce materially new short-circuit levels, but only refine already-sampled regions.
@@
Based on the converged Monte Carlo set, the resulting distributions of the total wind-farm penetration and of the maximum turbine-to-turbine difference across all Monte Carlo realizations were illustrated in Fig. 8. The final Monte Carlo set encompass the 15,900 fault cases analyzed in this study.
@@ V Assessment of the Methodology
The evaluation employed the simulation scenarios described in Section 6, enabling a consistent comparison of fault-location accuracy between the best conventional approaches and the proposed method. Following the recommendations in [DAVI2025MULTI], the TAKZ [TAKAGI1982] algorithm was adopted for single-line-to-ground (SLG) faults, the enhanced TAKZ_New [MOISESLOC1] formulation for double-line-to-ground (DLG) faults, the TAKN [SEL2018] method for line-to-line (LL) faults, and the Reactance [CAPAR2014] approach for three-phase (3P) faults. The overall performance across these four major fault classes is summarized in Fig. 9.
@@
Complementing this qualitative insight, Table IV quantifies the average location error for each fault type, with and without the proposed compensation. These results corroborate the trends observed in Fig. 9, i.e., average errors for SLG and DLG faults drop from 0.584% and 0.386% to 0.065% and 0.095%, respectively (improvements that exceed 75%). LL and 3P faults also benefit, with error reductions of approximately 46% and 16%, respectively. These findings confirm that the compensation term significantly mitigates the distortions caused by IBRs, especially for faults that are not dominated by positive-sequence contributions.
@@
Finally, Fig. 11 compares the error distribution between faults occurring in the primary line (upstream of any inverter) and the secondary line, where IBR injections are present between the IED and the fault. As expected, in the primary line, compensated and uncompensated errors were nearly identical as no IBRs intervene in the loop. However, for the secondary line, the uncompensated error was significantly larger, reflecting the distortion introduced by unaccounted inverter injections. The compensated results in the secondary line closely match those of the primary line, demonstrating that the methodology effectively attenuates the influence of IBRs when their position lies between the IED and the fault.
@@
Taken together, the results shown in these figures and tables demonstrate that the proposed compensation methodology offers substantial and consistent gains in fault location accuracy across fault types, system conditions, and topological configurations. Its ability to mitigate the distortive effects of inverter injections, particularly in complex wind farm collector systems, suggests that it can serve as a robust enhancement to conventional one-terminal phasor-based locators.
@@ VI Conclusion
This paper highlighted a fundamental limitation in traditional one-terminal phasor-based fault location methods when applied to wind farm collector networks, namely their reduced accuracy in the presence of IBRs located downstream from the fault point. To address this challenge, a simple and general compensation methodology was proposed to correct the loop quantities affected by IBR current injections. The framework maintains the structural simplicity of classical approaches while significantly improving fault location accuracy in IBR-dominated environments.
The methodology was validated through an extensive set of simulations conducted in PSCAD using a realistic onshore wind farm model comprising 27 full-converter turbines. More than 15,000 fault cases were evaluated, spanning variations in fault type, resistance, inception angle, distance, and wind penetration level. For each scenario, the proposed method was applied in conjunction with the best-performing phasor-based locator for that specific fault condition.
Quantitative results demonstrated that the proposed compensation strategy reduces both average and maximum fault location errors. Improvements exceeding 90% were observed for SLG and DLG faults. Significant error reductions were also achieved for line-to-line and three-phase faults, despite their partial reliance on sequence components affected by IBR dynamics. Furthermore, the methodology was demonstrated to mitigate the sensitivity of fault location error to variations in wind power penetration and to equalize performance across feeder segments, regardless of whether IBRs are present between the fault and the measurement point.
In conclusion, this addresses a previously underexplored challenge in fault location for wind farm collector systems with high IBR integration. The proposed compensation framework provides a low-complexity, method-agnostic enhancement to traditional phasor-based schemes, thereby restoring fault location accuracy in scenarios where conventional methods exhibit low accuracy. As renewable penetration increases, such techniques become essential for ensuring fast and reliable system restoration following faults.

### v2

@@ Abstract
The increasing integration of Inverter-Based Resources (IBRs) is reshaping fault current characteristics, presenting significant challenges to traditional protection and fault location methods. This paper addresses a key limitation in fault location within wind farm collector networks, i.e., one-terminal phasor-based methods become inaccurate when IBRs are electrically located downstream from the fault. In such cases, the voltage drop caused by IBR fault current injections is not captured by the Intelligent Electronic Device (IED), resulting in a systematic overestimation of fault distance. To mitigate this issue, a general compensation framework was proposed by augmenting classical loop formulations with a distance-dependent voltage correction term. The methodology was derived analytically using a sequence-domain representation and generalized to multiple fault types through a unified notation. It maintains the simplicity and interpretability of conventional approaches and can be implemented using only local measurements. The method was evaluated through EMT simulations in PSCAD using a realistic wind farm model. Results show significant improvements in location accuracy, with average and maximum errors notably reduced, especially for ground-involved faults where reductions exceed 90%. Furthermore, the compensation eliminates sensitivity to wind penetration levels and ensures uniform performance across feeders, positioning the method as a practical solution for modern renewable-dominated grids.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Most fault-location methods are categorized into four groups: phasor-based, voltage-sag, traveling-wave, and data-driven. Among these, phasor-based methods are often preferred because they are straightforward to implement, interpretable, and compatible with existing measurement infrastructures [DAS2014].
Within phasor-based methods, single-ended schemes rely exclusively on local measurements to estimate the fault distance. Common methods include the impedance-based formulation [ZIEGLER2011], reactance-based estimators [CAPAR2014], and the family of Takagi-type solutions [TAKAGI1982, DAS2014, SEL2018]. Two-ended schemes extend this approach by incorporating synchronized measurements at both line terminals, which requires a communication channel; notable contributions include [GIRGIS1992, JJ1990, PR2011, HE2011]. Despite this variety, phasor-based methodologies are grounded in standard short-circuit assumptions associated with synchronous-machine behavior.
@@
In transmission systems with high IBR penetration, two complementary research directions have emerged. For instance, in [Davi2023IBRMultiMethod] the authors assessed ten impedance-based fault-location schemes, comprising both one- and two-terminal methods, using several PSCAD simulations covering a wide range of IBR control strategies and short-circuit levels. Based on these results, they proposed a multi-method selection scheme that significantly reduces the location error when no communication channel is available.
The second direction focuses on sequence-aware single-ended formulations tailored to the converter behavior. In [Mobashsher2024LocalIBR], the authors developed a sequence-network model for IBRs showing that the apparent negative-sequence impedance became effectively infinite when grid codes did not mandate negative-sequence injection, and designed a single-ended locator that employed negative- and zero-sequence polarizing currents, incorporated remote-infeed and line-charging effects, and maintained accuracy across wide ranges of fault resistance and grid-code settings.
@@
In order to address these limitations, this paper presents an enhanced single-ended, phasor-based fault-location method specifically designed for wind-farm collector systems. The key contribution is a transparent modification of the classical loop equation that explicitly incorporates the voltage drops caused by intermediate IBR-injected currents. By doing so, the proposed approach enables high-accuracy fault location in renewable-rich environments while preserving the simplicity and local-measurement dependence of conventional schemes.
@@ III-A Single-Line-to-Ground Fault Derivation
Together with the positive-sequence relation in Equation (9), these expressions fully specify the sequence-domain fault voltages ($V_{F}^{(1)},V_{F}^{(2)},V_{F}^{(0)}$). For other fault types, the per-sequence voltage relations themselves remain unchanged; what differs is the interconnection of the sequence networks at the fault. For example, in an SLG fault, the positive-, negative-, and zero-sequence networks are connected in series, whereas in a line-to-line fault, the positive- and negative-sequence networks are in series while the zero-sequence branch is open.
@@ III-B Extension to Other Fault Types
As the compensation term $V_{\text{comp}}(d)$ depends explicitly on the unknown distance $d$, its evaluation requires an iterative procedure. The estimation starts by obtaining an initial value of distance from (16) with $V_{\text{comp}}(d)=0$. This preliminary estimate is then substituted into (17) to update the compensation term, yielding a new value of $V_{\text{comp}}(d)$. The updated compensation is subsequently reinserted into (16) to produce a refined estimate of $d$. This fixed-point iteration is repeated until the difference between successive distance estimates falls below a prescribed tolerance, ensuring convergence.
@@ III-C Practical Considerations and Iterative Estimation Procedure
Assuming that: (i) all inverter generations are approximately equal during and previous the fault; and (ii) their positive-sequence injections maintain magnitude and phase angle close to pre-fault values, those assumptions can lead to an approximated form of Equation (17) as shown next:
@@ IV Test System
All WTGs were modeled using full-converter technology with a grid-following (GFL) control architecture [TREMBLAY2013, MILLER2003]. The GFL inverter synchronizes with the grid via a synchronous-reference-frame phase-locked loop (PLL) [Se2000] and regulates the current through cascaded control loops. The inner current-control loop operates in the $dq$ frame and enforces the commanded current using PI controllers tuned via internal model control [Harne1998]. The outer loop manages active and reactive power set-points, with a default operation at a unity power factor during steady-state. The current magnitude was constrained to 1.1 $1\text{\,}\mathrm{p}\,\mathrm{u}\,$, and fault ride-through behavior was incorporated such that when the point of common coupling (PCC) voltage drops below $0.85\text{\,}\mathrm{p}\,\mathrm{u}\,$, the control system injects positive-sequence reactive current [Dadjo2023], as mandated by the Brazilian grid code [ONS2023].
@@ IV-A Monte Carlo Generation of Wind-Penetration Scenarios
The accuracy of the proposed fault-location methodology was influenced not only by the fault location and resistance, but also by the spatial distribution of wind-power injections along the collector system. The position of each wind turbine and its local operating point affect the pre-fault current sharing and, consequently, the short-circuit level seen at the relay. In real installations, however, only the aggregated measurements at the IED bus are typically available, and the individual turbine currents are not recorded. In this work, each IBR contribution was therefore approximated as the pre-fault current at the IED divided by the number of WTGs (see Section III-C), implicitly assuming identical power output among turbines. Any dispersion in turbine output around this average may introduce additional location errors in the proposed methodology. To obtain a realistic and systematic assessment of this sensitivity, a Monte Carlo [rubinstein_simulation_2016] model was adopted to generate a large set of penetration scenarios that explicitly accounts for turbine-to-turbine variability.
@@
Thus, across scenarios, the wind farm operates between $0\%$ and $100\%$ of rated power. The deviations were independently drawn from a symmetric uniform law as:
@@
Fig. 7 shows the evolution of $\varepsilon_{99}(n)$ as a function of the number of scenarios, together with the convergence threshold. As the number of simulations increases, the nearest-neighbor distance decreases approximately monotonically, indicating progressively finer resolution of the short-circuit current distribution. The resolution criterion was satisfied only when the Monte Carlo set approaches its final size, implying that additional simulations beyond this point would not introduce materially new short-circuit levels, but only refine already-sampled regions.
@@
Based on the converged Monte Carlo set, the resulting distributions of the total wind-farm penetration and of the maximum turbine-to-turbine difference across all Monte Carlo realizations were illustrated in Fig. 8. The final Monte Carlo set encompass the 17,098 fault cases analyzed in this study.
@@ V Assessment of the Methodology
The evaluation employed the simulation scenarios described in Section 6, enabling a consistent comparison of fault-location accuracy between the best conventional approaches and the proposed method. Following the recommendations in [DAVI2025MULTI], the TAKZ [TAKAGI1982] algorithm was adopted for SLG faults, the enhanced TAKZ_New [MOISESLOC1] formulation for double-line-to-ground (DLG) faults, the TAKN [SEL2018] method for line-to-line (LL) faults, and the Reactance [CAPAR2014] approach for three-phase (3P) faults. The overall performance across these four major fault classes is summarized in Fig. 9.
@@
Complementing this qualitative insight, Table IV quantifies the average location error for each fault type, with and without the proposed compensation. These results corroborate the trends observed in Fig. 9, i.e., average errors for SLG and DLG faults drop from 0.584% and 0.386% to 0.065% and 0.095%, respectively (improvements that exceed 75%). LL and 3P faults also benefit, with error reductions of approximately 46% and 16%, respectively. These findings confirm that the compensation term significantly mitigates the IBR-induced deviations, especially for faults that are not dominated by positive-sequence contributions.
@@
Finally, Fig. 11 compares the error distribution between faults occurring in the primary line (upstream of any inverter) and the secondary line, where IBR injections are present between the IED and the fault. As expected, in the primary line, compensated and uncompensated errors were nearly identical as no IBRs intervene in the loop. However, for the secondary line, the uncompensated error was significantly larger, reflecting the deviations introduced by unaccounted inverter injections. The compensated results in the secondary line closely match those of the primary line, demonstrating that the methodology effectively attenuates the influence of IBRs when their position lies between the IED and the fault.
@@
Taken together, the results shown in these figures and tables demonstrate that the proposed compensation methodology offers substantial and consistent gains in fault location accuracy across fault types, system conditions, and topological configurations. Its ability to mitigate the IBR-induced deviations, particularly in complex wind farm collector systems, suggests that it can serve as a robust enhancement to conventional one-terminal phasor-based locators.
@@ VI Conclusion
This paper highlighted a fundamental limitation in traditional one-terminal phasor-based fault location methods when applied to wind farm collector networks, namely their reduced accuracy in the presence of IBRs located downstream from the fault point. To mitigate this effect, a simple compensation framework was proposed to correct loop quantities distorted by IBR current injections, preserving the structural simplicity of classical schemes while enhancing their accuracy under high IBR penetration. The methodology was assessed in PSCAD using a realistic onshore wind farm with 27 full-converter turbines and more than 15,000 fault scenarios spanning different fault types, resistances, inception angles, locations, and wind penetration levels.
Quantitative results showed substantial reductions in both average and maximum fault-location errors, exceeding 75% for SLG and DLG faults, and significant improvements for LL and 3P faults. The proposed approach also reduced the sensitivity of the locator to variations in wind penetration and helped equalize performance along the feeder, irrespective of IBR placement. Overall, the compensation framework provides a low-complexity, method-agnostic enhancement that restores the effectiveness of phasor-based fault-location schemes in wind-farm collector systems with high IBR integration.
Acknowledgments
The authors would like to acknowledge the support of the National Council for Scientific and Technological Development under Grant No. 307982/2022-0. They also gratefully acknowledge the Research Centre for Greenhouse Gas Innovation at the University of São Paulo, funded by the São Paulo Research Foundation under Grant No. 2020/15230-5 and by TotalEnergies, as well as the strategic support provided by the Brazilian National Oil, Natural Gas and Biofuels Agency through the R&D levy regulation.